OpenStudy (anonymous):

what is the slope of 1/3, (-3, 0)

OpenStudy (doc.brown):

Are you cutting your line going -x in 1/3?

OpenStudy (anonymous):

wat is ur exact question?????

OpenStudy (nirmalnema):

u can use one point form... (y-y1)=slope (x-x1) y=1/3 (x+3) 3y=x+3 x-3y+3=0